OpenStudy (anonymous):

Solve? 3^(4x-7)= 4^(2x+3)

OpenStudy (anonymous):

Try taking the ln of each side, and use logarithmic properties to simplify for x?

OpenStudy (anonymous):

x is around 7.30591

OpenStudy (athe):

\[\frac{9^{2x}}{3^7}=4^{2x}4^3\] \[\frac{9^{2x}}{4^{2x}}=3^74^3\] \[(\frac{9}{4})^{2x}=3^74^3\] \[x=\frac{1}{2} \log_{9/4} (3^7 4^3)\]

sam (.sam.):

3^(4x-7)=4^(2x+3) Take the natural logarithm of both sides of the equation to remove the variable from the exponent. ln(3^(4x-7))=ln(4^(2x+3)) The left-hand side of the equation is equal to the exponent of the logarithm argument because the base of the logarithm equals the base of the argument. 4xln(3)-7ln(3)=ln(4^(2x+3)) The exponent of a factor inside a logarithm can be expanded to the front of the expression using the third law of logarithms. The third law of logarithms states that the logarithm of a power of x is equal to the exponent of that power times the logarithm of x (e.g. log^b(x^(n))=nlog^b(x)). 4xln(3)-7ln(3)=((2x+3)ln(4)) Multiply ln(4) by each term inside the parentheses. 4xln(3)-7ln(3)=(2xln(4)+3ln(4)) Since 2xln(4) contains the variable to solve for, move it to the left-hand side of the equation by subtracting 2xln(4) from both sides. 4xln(3)-7ln(3)-2xln(4)=3ln(4) Factor out the GCF of 2x from each term in the polynomial. 2x(2ln(3))+2x(-ln(4))=7ln(3)+3ln(4) Factor out the GCF of 2x from 4xln(3)-2xln(4). 2x(2ln(3)-ln(4))=7ln(3)+3ln(4) Divide each term in the equation by (2ln(3)-ln(4)). (2x(2ln(3)-ln(4)))/(2ln(3)-ln(4))=(7ln(3))/(2ln(3)-ln(4))+(3ln(4))/(2ln(3)-ln(4)) Simplify the left-hand side of the equation by canceling the common factors. 2x=(7ln(3))/(2ln(3)-ln(4))+(3ln(4))/(2ln(3)-ln(4)) Simplify the right-hand side of the equation by simplifying each term. 2x=(7ln(3)+3ln(4))/(2ln(3)-ln(4)) Divide each term in the equation by 2. (2x)/(2)=(7ln(3)+3ln(4))/(2ln(3)-ln(4))/(2) Simplify the left-hand side of the equation by canceling the common factors. x=(7ln(3)+3ln(4))/(2ln(3)-ln(4))/(2) Simplify the right-hand side of the equation by simplifying each term. x=(7ln(3)+3ln(4))/(2(2ln(3)-ln(4))) x=0.6166

OpenStudy (anonymous):

take log both sides, think someone else asked this aswell: log3^(4x-7)=log4^(2x+3) 4x-7/2x+3=log4/log3 4x-7/2x+3=1.26 (2x+3)(1.26)=4x-7 2.52x+3.78=4x-7 3.78+7=4x-2.52x 10.78=1.48x 10.78/1.48=x 7.28=x